Minutes BWS Upgrade: Electro-Mechanical Design Meeting #31 -- 12th June 2018 --

Participants: R. Veness, F. Roncarolo, J. Emery, P. Andersson, D. Gudkov, J.L. Sirvent

Agenda (https://indico.cern.ch/event/729280/)

1. Series Production Production (D.Gudkov)

  • All parts delivered except:
    • Tanks will all be delivered at the same time, end of November 2018.
    • Delivery dates - PS in Feb 2019-June, Booster july, SPS September
  • Visit from Emilien in ~3 weeks
  • Prototype tank from vacuum project received (for the PSB) at CERN. Should be complete mid July.
  • Metallic discus - 30 blanks ordered RSA905 with LTultra. Not yet placed order for engraving.
  • Shaft is compatible with all discs. Some difference in inertia. Series shaft has longer thread, so would be better for glass disc.
  • Series should be ready for installation in September.
  • 3/4 stepped chambers are ready.
  • Yvon Boyer managed to machine the key slots properly in the shafts, so will complete the series with them
  • Viewports - Can order 30 viewports from Vac Com with 4-5 weeks delivery. Agree for the moment to only put on 2 viewports per instrument.
  • Assembly area is ready. All kinematic units ready for end of September.
  • Photographer coming tomorrow afternoon.
  • Would like to make some tests of kinematic unit in the lab - perhaps from this summer, so that we can make some burn-in cycles.
  • There si a d-space system with rack and wheels being assembled. This is probably the only way to make nominal cycles.
  • Will come back in the next meeting and report on what is practical. Could imagine to do this in batches.

2. Electronics Progress (J.Emery):

  • There is enough hardware that could be assembled for a test. Missing software and firmware. However, iterations are in-progress with design office (eg for the Mezzanine).
  • There was an internal review last week for the three cards. Made some suggestions for EM compatibility etc, and was a positive process.
  • Software section ‘on hold’ due to resolver issue. New staff (Lioc Fosse) just started, but not yet operational. Firmware now being worked-on with Status with new VME board.
  • Started on architecture with Ana G.
  • Risks - only test in one machine so different EM environment. Different VMC architecture. Effect of the beam on the acquisition chain (but this will be done ASAP).
  • Confident that performance test and calibration bench with be enough.
  • SPS cable run of ~200m for 519, this is less than current 290m in SPS. Currently 185m installed in longest run. Problem with loss of current in the cable run. They have a fix that would improve this.
  • Ray ask how much margin we have in the system. No clear answer for now. Will be able to answer in July when we make tests in B.867.
  • PMT base - PCBs should be delivered tomorrow, but housing perhaps not before July.
  • Resolver filter (update from last meeting), now 600 scans with beam. 3 to investigate as some issues.
  • Jonathan proposes 1280 CHF for 2x digital meters that will allow a logging of the wire continuity at all times.
  • Dmitry will send revised code splitting including ESS series.

3. Test results from PS (J.L.Sirvent)

  • 631 scans in total in the PS.
    • LHC probe, ~6e9 protons
    • LHCINDIV -singel buch with 0.8x10e11 Protons
    • TOF with ~7e12 protons
  • No sign of disc drift
  • Jitter 205 uS comparing disc with position.
  • Possibly some drift of the glass disc - to be confirmed.

4. Holidays (R.Veness)

  • Jose spain 3 weeks end-july then back ~15 august
  • Dmitry next wednesday for 2 weeks upto 4 july then 15-25 august
  • Federico last week of july then 15 august for some days
  • Jonathan june for training then last 2 weeks in august
  • Patrick 2 first weeks in August

5. TS1 Interventions (R.Veness)

5.1. PSB
  • Agree we change the scanner if possible, if not the change wires on the installed scanner.
  • Need to check on the solution for re-fixing the Kapton wire in place.
5.2. SPS
  • 521 intervention is ready for linear scanner. Wire is OK.
  • Do we update the software in the SPS?
5.3 PS
  • Nothing to do in the shutdown
TS2 is 17-18 September
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2018-06-13 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    BWSUpgrade All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback